Hello! I am taking finite math and the teacher never showed us a problem like this ( Only cards, dice and urns) and expects us to know how to do it. please help! The Question:
Suppose 5 men out of 100 and 25 woman out of 10,000 are colorblind. A colorblind person is chosen at random. What is the probability of his being male? (Assume males and females to be in equal numbers) I don't understand what the sample space is. Is it 10,000 or 10,000 twice? Please just show me how you solve this problem with the answer so that I can understand what type of probability problem this is and be able to solve any like it. Thanks so much!! Stacy

Answer by ewatrrr(24785)

p(M cb) = 5/100 = .05, p(F cb) = 25/10,000 = .0025
......
P(A|B) = P(A and B)/P(B)
P(M | cb) = .05/.0525
P(cb)=(Adding .05 + .0025)